Nicole:

The circumference of a circle with the diameter of 3 cm is 9.42 cm. What is the ratio of the circumference to the diameter?

dude:

I think you divide circumference by diameter, I haven't done this in the longest so I am rusty in this \(\large{\frac{9.42}{3}}\)

Nicole:

ok

Nicole:

How would you use a model to determine the approximate formula for the area of a circle with a diameter of 8 cm, and connect the model to the actual formula?

Nicole:

What about that one?

satellite73:

yikes! the ratio of the circumference of the circle to the diameter is always \(\pi\) that is a good definition of \(\pi\) one wonders what kind of math teacher writes questions like these?

sillybilly123:

so if we hypothesise that the ratio of the circumference (C) of the circle to the diameter (D) is actually always some constant \(\xi\): ie \(\dfrac{C}{ D} = \xi\), ....then we set about finding the value of \(\xi\) So how do we do that? This is actually, IMHO, a good teaching strategy. If you think it's obvious that \(\xi = \text{whatever}\), then go and prove it !! And then prove that it is \(\xi\) is constant